We estimate a linear probability model with PTSD as our dependent variable and the number of years on the force as our only independent variable. We get the following results:


Coefficientsa Model Unstandardized Coefficients Standardized Coefficients t Sig. B Std. Error Beta (Constant) .702 .064 10.969 .000 Number of Years on Force -.043 .006 -.609 -7.167 .000 R2 = .37 with these results, answer the following questions:


a. Write out the full regression equation.
b. Interpret the constant.
c. Interpret the slope coefficient.
d. Using an alpha of .05, what is your decision about the null hypothesis: byrsontheforce = 0? Explain.
e. Interpret the value of R2.
f. What is the predicted probability of having PTSD for an officer who has 5 years on the force?
g. What is the predicted probability of having PTSD for an officer who has 15 years on the force?
Answer #1

the full regression equation as per your results shown here is

a) PTSD = 0.702 - *0.043( Number of Years on Force )

b) Interpret the constant: When the number of years on Force is zero which means for a person who doesn't work in Force the PTSD = 0.702

c) Interpret the slope coefficient: If the number of years on Force is increased by one unit there will be 0.043 units decrease in PTSD.

d) null hypothesis: byrsontheforce = 0

the p-value of the coefficient of number of years on Force is zero which is less than 0.05 and concludes that we have to reject null hypothesis and there exists a significant linear relationship between the variables.

e) Interpret the value of R2 :- R2 = 0.37 which says that the model explains only 37% of the variability of the PTSD with number of years on Force.

f) 5 years on the force

PTSD = 0.702 - 0.043*5 = 0.487

g) 1 5 years on the force

PTSD = 0.702 - 0.043*15 = 0.057